Last night, I plugged the ETMX suspension coils back into the satellite box. Tonight, we turned on the damping loops for ETMX. Rana centered the Oplev so we can use that as an additional diagnostic to see if the optic gets kicked around overnight. We will re-assess the situation tomorrow.
Sometime earlier today, Lydia noticed that the +/- 5V Sorensens at the X end were not displaying their nominal voltage/current values (as per the stickers on them). She corrected this.